the average height of randomly chosen 45 female students was 155cm. with standard deviation of 3.5 cm find the 95% confidence interval


1
Expert's answer
2022-05-01T17:26:52-0400

The formula to calculate a confidence interval for a population mean is as follows:

"CI=\bar{x}\pm z\cdot\cfrac{s}{\sqrt{n}},"

where:

  • "\bar{x}=155" : sample mean
  • zthe chosen z-value, for a 95% confidence interval z = 1.960
  • "s=3.5" sample standard deviation
  • n = 45 sample size.

So,

"CI=155 \pm 1.960\cdot\cfrac{3.5}{\sqrt{45}}=\\=
155\pm1.02=(153.98, 156.02)."